# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

### v2

@@ 4 Related Work
We identify three primary categories on retrieval-oriented question answering (QA) datasets based on the knowledge sources they utilize: unstructured, structured, and semi-structured knowledge bases. We discuss the differences and limitations in the previous works that we seek to address.
Unstructured QA Datasets. The line of work explores retrieving answers from a single document [33] or multiple document sources [42, 47, 22, 44, 10]. For example, SQuAD [33] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context. For multi-document retrieval, HotpotQA [47] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task. TriviaQA [22] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources. Moreover, some works [30, 25] leverage generated results by search engines as ground truth answers or information sources for retrieval. However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries. In contrast, STaRK contains queries demanding multi-hop relational reasoning to challenge model’s ability of handling structured information.
Structured QA Datasets. Within this category, the datasets challenge models to retrieve answers from structured knowledge bases, either knowledge graphs [3, 51, 40, 4, 5, 11, 14] or tabular data [53, 52]. For example, ComplexWebQuestions [40] challenge models’ abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs. GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language models a way to handle KBQA. For datasets with tabular data, WikiSQL [53] and Spider [52] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables. While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information. STaRK extends this line of work by integrating textual information into knowledge bases and queries, resulting in more comprehensive and nuanced responses with extensive insights drawn from abundant textual data.
Semi-Structured QA Datasets. Another line of work focuses on semi-structured QA with tabular and textual data. WikiTableQuestions [32] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets. TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types. TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data. However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information that naturally exists between entities. Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate. Therefore, STARK goes beyond these limitations by integrating both relational and textual data in a more cohesive manner, allowing for seamless navigation and retrieval across complex semi-structured knowledge bases. This integration supports richer relational reasoning and more comprehensive textual data handling, making STARK a more effective benchmark for evaluating the capabilities of retrieval systems in scenarios that closely mirror real-world information needs.
5 Conclusion
@@
We introduce STaRK, the first benchmark designed to thoroughly evaluate the capability of retrieval systems driven by LLMs in handling semi-structured knowledge bases (SKBs). This benchmark features a diverse set of queries that are semi-structured and natural-sounding, requiring context-specific reasoning across various domains, thereby setting a new standard for assessing retrieval systems in the context of SKBs. We utilize public datasets to construct three SKBs and develop an automated, general pipeline to simulate user queries that mimic real-world scenarios. Moreover, we augment our datasets with human-generated queries and provide reference to the synthesized queries. Our extensive experiments on STaRK reveal significant challenges for current models in handling both textual and relational information effectively and flexibly. Overall, STaRK offers valuable opportunities for future research to advance the field of complex and multimodal retrieval systems, where reducing retrieval latency and incorporating strong reasoning ability into the retrieval process are identified as two prospective future directions.
@@ 6 Acknowledgement
We thank group members group members in Leskovec and Zou labs for providing valuable suggestions and conducting benchmark construction. We express our gratitude to the following individuals for their assistance in generating the human-generated queries (ordered by last name):
@@
We also gratefully acknowledge the support of DARPA under Nos. N660011924033 (MCS); NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
Appendix A Mathematical Definitions of Shannon Entropy and Type-Token Ratio
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$
$\text{TTR}=\frac{V}{N}$
Appendix B Benchmark details
@@
B.1 STaRK-Amazon
Basic query templates:
metapath Query template
(brand → product) "Can you list the products made by <brand>?"
(product → product) "Which products are similar to <product>?"
The final relational template can be composed of multiple basic templates. For example, ‘(color → product ← brand)’ represents one such combination.
B.2 STaRK-MAG
Relational query templates:
metapath multi-hop query template
(author → paper) "Can you list the papers authored by <author>?"
(paper → paper) "Which papers have been cited by the paper <paper>?"
B.3 STaRK-Prime
Number of entities in each type.
@@ B.3 STaRK-Prime
Relational query templates.
@@
    "What drug has a synergistic interaction with the drug that has <effect/phenotype> as a side effect?",
@@
    "What biological process has the common interactino pattern with gene or proteins as <biological_process>?",
@@
    "Find drugs that has a synergistic interaction with <drug> and both are indicated for the same disease.",
@@
    "Find gene/protein that can interect with <gene/protein> and both are associated with the same disease.",
@@
    "Find gene/protein that can interect with <gene/protein> and both are associated with the same effect/phenotype."
@@
Appendix C Prompts and LLM versions for Query Synthesization
We summarize the LLM versions in Table 11. We chose these models based on their cost, how accurate they are, and whether they were the latest model during different parts of the project. While we used different LLMs, we checked each step separately to make sure the good quality in our benchmark datasets.
@@ Appendix C Prompts and LLM versions for Query Synthesization
C.1 Extracting textual requirements
@@
C.2 Combining relational and textual requirements
@@
C.3 Filtering additional answers
@@
Please provide only the JSON in your response. Response:"
Appendix D Instructions for Generating Queries
For the process of generating queries by human, the participants were given a list of entity IDs that we randomly sampled from the entire entity set. Then, they were asked to follow the following instructions with the support of our built interactive demo at https://huggingface.co/spaces/snap-stanford/SKB-Explorer. {mdframed}[backgroundcolor=gray!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt] Task:
Note:
1) Please do not leak the name of the entity in the query.
2) You can skip some queries if you think the knowledge involved is hard to understand.
3) Feel free to be creative with content of your queries, you can also include additional context. There is NO restriction on how you express the queries. After that, we filtering the ground truth answers manually by human validation.

### v3

@@ 4 Related Work
Unstructured QA Datasets. This research domain consists of methods for retrieving answers from unstructured text, either from a single document [39] or multiple documents [50, 55, 25, 52, 13]. For instance, SQuAD [39] is designed for answer extraction within a specific document, while approaches like HotpotQA [55] and TriviaQA [25] extend to multi-document contexts. Additionally, some studies utilize search engine outputs as a basis or supplementary data for question answering [35, 29]. However, unstructured QA datasets often lack the depth of relational reasoning commonly required in answering complex user queries. In contrast, STaRK contains queries demanding multi-hop relational reasoning to challenge model’s ability of handling structured information.
Structured QA Datasets. These datasets challenge models to derive answers from structured sources such as knowledge graphs [6, 59, 48, 7, 8, 14, 17] or tabular data [61, 60]. ComplexWebQuestions [48] and GraphQA [17] propose challenges in interpreting complex queries and textualizing graph structures in KBQA, respectively. For tabular data, WikiSQL [61] focuses on translating queries to SQL for single-table databases, whereas Spider [60] tackles multi-table scenarios. Despite the emphasis on relational data, the restriction to predefined entities and relationships limits the scope of queries. STaRK integrates textual content within structured frameworks to enhance the depth and breadth of information retrieval, promoting richer and more nuanced understanding from extensive textual data.
Semi-Structured QA Datasets. This category merges tabular and textual data, presenting challenges in semi-structured data comprehension. WikiTableQuestions [38] stresses the integration of table structures with textual elements. TabFact [10], HybridQA [11], and TabMCQ [23] extend this by combining validation of textual statements with tabular reasoning. However, datasets leveraging tables as structured frameworks often lack in depicting the rich relational dynamics among entities. Moreover, prior efforts to link textual and tabular information via external sources have led to cumbersome data constructs. Addressing these challenges, STaRK enhances integration, allowing for flexible navigation and advanced retrieval within complex semi-structured knowledge bases, and facilitating more effective relational reasoning and text handling.
5 Conclusion and Future Work
@@
We introduce STaRK, the first benchmark to thoroughly evaluate LLM-driven retrieval systems for semi-structured knowledge bases (SKBs). Featuring diverse, natural-sounding queries that require context-specific reasoning across diverse domains, STaRK sets a new standard for assessing real-world retrieval systems. We contribute three large-scale retrieval datasets with human-generated queries and an automated pipeline to simulate realistic user queries. Our experiments on STaRK highlight significant challenges for current models in effectively handling textual and relational information. STaRK paves the way for future research to advance complex, multimodal retrieval systems, focusing on reducing retrieval latency and enhancing reasoning abiliites.
Our current SKBs are limited to textual and relational information. Future work should incorporate additional modalities such as images, videos, and speech to provide a more comprehensive information retrieval system. Despite our anonymization efforts, we acknowledge that privacy remains a potential concern when extending this work to other domains with real user data, which should be protected to ensure compliance with privacy regulations.
@@ 6 Acknowledgement
We thank group members in Leskovec and Zou labs for providing valuable suggestions and conducting benchmark construction. We express our gratitude to the following individuals for their assistance in generating the human-generated queries (ordered by last name):
@@
We gratefully acknowledge the support of DARPA under Nos. N660011924033 (MCS); NSF under Nos. OAC-1835598 (CINES), CCF-1918940 (Expeditions), DMS-2327709 (IHBEM); Stanford Data Applications Initiative, Wu Tsai Neurosciences Institute, Stanford Institute for Human-Centered AI, Chan Zuckerberg Initiative, Amazon, Genentech, GSK, Hitachi, SAP, and UCB.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] We clearly state our problem scope and highlight our main contribution compared to the existing works.
    2. (b)
      Did you describe the limitations of your work? [Yes] Please see the conclusion and future work section.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] Please see the conclusion and future work section.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Our website https://stark.stanford.edu/ and GitHub codebase https://github.com/snap-stanford/STaRK contains code, data, and experimental pipelines.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Please see the experiment setup in Section 3.1, where we explained all of the choices made. We also make the data splits public.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [N/A]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We provide the GPU device information and report the latency cost.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes] We mentioned them in our webpage.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] The resources are from exisiting public data that is open to access.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] The only data with potential personally identifiable information and offensive content is Amazon semi-structured dataset, which is already made anonymized by the public resources.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] . We invited volunteered participants who are acknowledged.
Appendix A Benchmark details
A.1 Semi-structured Knowledge Bases (SKBs)
We present the public sources that we used to construct the SKBs in the table below. We have adhered to the licenses of each public resource.
@@ A.1 Semi-structured Knowledge Bases (SKBs)
We build an interactive platform to inspect the data of all three SKBs at https://stark.stanford.edu/skb_explorer.html. We introduce more detailed data statistics below.
Amazon SKB. In total, it comprises around 1.0M entities (product entities: 0.9M, brand entities: 0.1M, category entities: 1.4k, color entities: 1.7k) and 9.4M relations (also_bought: 2.8M, also_viewed: 1.9M, has_brand: 1.7M, has_category: 2.3M, has_color: 0.6M).
MAG SKB. This SKB contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 8.7K, field_of_study: 59.5k) and 39.8M relations under four relation types (author_writes_paper: 13.5M, paper_has_field_of_study: 14.5M, paper_cites_paper: 9.7M, author_affiliated_with_institution: 2.0M).
Prime SKB. The entity count in our knowledge base is approximately 129.3K, with around 8.1M relations. The numbers of entities in each type are listed below:
@@
A.2 STaRK-Amazon
Relational query templates. These are the basic relational templates on STaRK-Amazon. Note that the final relational template can be composed of multiple basic templates. For example, ‘(color → product ← brand)’ represents a relational template combined from two basic relational templates.
metapath Query template
(brand → product) "Can you list the products made by <brand>?"
(product → product) "Which products are similar to <product>?"
A.3 STaRK-MAG
Relational query templates. We constructed seven relational templates below:
metapath multi-hop query template
(author → paper) "Can you list the papers authored by <author>?"
(paper → paper) "Which papers have been cited by the paper <paper>?"
A.4 STaRK-Prime
Relational query templates. For synthesized queries, we listed 28 multi-hop templates designed by experts to cover various relation types and ensure their practical relevance.
For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments.
{mdframed}
[backgroundcolor=gray!5, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt]
@@ A.4 STaRK-Prime
    "What drug has a synergistic interaction with the drug that has <effect/phenotype>
    as a side effect?",
@@
    "What biological process has the common interactino pattern with gene or proteins as
    <biological_process>?",
@@
    "Find drugs that has a synergistic interaction with <drug> and both are indicated
    for the same disease.",
@@
    "Find gene/protein that can interect with <gene/protein> and both are associated
    with the same disease.",
@@
    "Find gene/protein that can interect with <gene/protein> and both are associated
    with the same effect/phenotype."
@@
Appendix B Mathematical Definitions of Shannon Entropy and Type-Token Ratio
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log p(x_{i})$
$\text{TTR}=\frac{V}{N}$
Appendix C Instructions for Generating Queries
For the process of generating queries by human, the participants were given a list of entity IDs that we randomly sampled from the entire entity set. Then, they were asked to follow the following instructions with the support of our built interactive platform at https://stark.stanford.edu/skb_explorer.html.
{mdframed}
[backgroundcolor=gray!10, linecolor=black, linewidth=2pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=3pt, innerbottommargin=10pt] Task:
Note:
1) Please do not leak the name of the entity in the query.
2) You can skip some entity IDs if you think the knowledge involved is hard to understand.
3) Feel free to be creative with content of your queries, you can also include additional context. There is NO restriction on how you express the queries. After collecting the queries, we filtering the ground truth answers manually by human validation.
Appendix D Experiments
D.1 More Experimental Results
Table 10: Positive/Non-negative rates (%) from human evaluation.
Naturalness Diversity Practicality
STaRK-Amazon 73.6 / 89.5 68.4 / 89.5 89.5 / 94.7
STaRK-MAG 94.7 / 100 73.7 / 84.2 68.4 / 84.2
STaRK-Prime 67.8 / 92.8 71.4 / 82.1 71.4 / 89.3
Average 78.7 / 94.1 71.0 / 85.3 76.4 / 89.4
Appendix E Prompts and LLM versions for Query Synthesization
We summarize the LLM versions in Table 11. We chose these models based on a joint consideration of their cost, how accurate they are, and whether they were the latest model during different phases of the project. While we used different LLMs, we checked each step separately to make sure the good quality in our benchmark datasets.
@@ Appendix E Prompts and LLM versions for Query Synthesization
E.1 Extracting textual requirements
@@
E.2 Combining relational and textual requirements
@@
E.3 Filtering additional answers
@@
Please provide only the JSON in your response. Response: